OpenStudy (anonymous):

Find an exact value. cos 165°

Nnesha (nnesha):

whats two number you can add to get 165? and those two number suppose to be on the unit circle

Nnesha (nnesha):

then use formula \[\large\rm cos(a+b)=\cos a \times \cos b - \sin a \times \sin b\]

OpenStudy (anonymous):

90 & 75 ?

Nnesha (nnesha):

well 75 is not on the unit circle :/

Nnesha (nnesha):



OpenStudy (anonymous):

120 + 45

Nnesha (nnesha):

30+135 40+120 225-60 for this u have to use cos(a-b)=cos a times cos b + sin a times sin b

Nnesha (nnesha):

perfect!

Nnesha (nnesha):

so \[\huge\rm cos(45+120)=\cos 45 \cos 120 - \sin 45 \sin 120\]

OpenStudy (anonymous):

okay, i see. what do i do from there?

Nnesha (nnesha):

so \[\large\rm cos(45+120)=\cos 45 \cos 120 - \sin 45 \sin 120\] now use the unit circle cor exact value of cos and sin cos 45 =? cos 120 =? sin 45=? sin 120 =? plugin their exact values

OpenStudy (anonymous):

cos 45 = (radical)2 / 2 cos 120 = -1/2 sin 45 = (radical) 2 /2 sin 120 = (radical) 3 /2 is that right?

Nnesha (nnesha):

\[\frac{ \sqrt{2} }{ 2 } \times \frac{ -1 }{ 2 } - \frac{ \sqrt{2} }{ 2 } \times \frac{ \sqrt{3} }{ 2}\] solve

Nnesha (nnesha):

yes that's right